Let X be normally distributed with mean μ = 2.9 and standard deviation σ = 1.5. Use Table 1.

a.

Find P(X > 6.5). (Round "z" value to 2 decimal places and final answer to 4 decimal places.)


  P(X > 6.5)


b.

Find P(5.5 ≤ X ≤ 7.5). (Round "z" value to 2 decimal places and final answer to 4 decimal places.)


  P(5.5 ≤ X ≤ 7.5)


c.

Find x such that P(X > x) = 0.0594. (Round "z" value and final answer to 3 decimal places.)


  x


d.

Find x such that P(xX ≤ 2.9) = 0.4066. (Negative value should be indicated by a minus sign. Round "z" value and final answer to 3 decimal places.)


  x
Answer #1

Given u= 2.9 and 0 = 1.5

a)- μ Ζ =- 6.5 – 2.9 1.5 3.6 -15 = 2.4

P(X > 6.5) = P(Z > 2.4) = 0.0082

b)ZO X1 - 5.5 - 2.9 1.5 2.6 15

X2-4 Z2 = 7.5 - 2.9 1.5 4.6 1.5 0

P(5.5 < X < 7.5) = P(1.73<Z <3.07) = 0.0404

c)Z = 1.56

r = 4 + Zo = 2.9+(1.56)1.5 = 5.24

d)2 = -1.32

I=+ Zo = 2.9+(-1.32)1.5 = 0.92